OpenStudy (anonymous):

1. What is the distance between P(12, 4) and Q(-8, 2)? 2. What is the distance between P(4, 1) and Q(12, -5)? 3. What is the distance between P(12, 11) and Q(5, 1)? 4. What is the distance between P(-3, -25) and Q(50, -2)? 5. Find the midpoint of if P (-3, 0) and Q(5, 12)? 6. Find the midpoint of if P (10, 2) and Q(8, 0)? 7. Find the midpoint of if P (-3, -4) and Q(5, -5)? 8. Find the midpoint of if P (10, 20) and Q(15, 20)?

OpenStudy (opcode):

Use the distance formula to find the distance between two points: \[\sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\] Use the midpoint formula to find the midpoint between two points: \[\dfrac{x_1+x_1}{2},\dfrac{y_1+y_2}{2}\]

OpenStudy (mathmale):

@silkdurag : would you please get started on the first problem, using the distance formula posted by Opcode, and showing your work? I'd bet you'd find the remaining problems relatively easy after you've finished the first one successfully.